# Locally homogeneous Axiom A flows II: geometric structures for Anosov subgroups

### v2

@@ 1.1. Dynamical $(G,X)$-structures
Here we give a short description of the homogeneous spaces studied in this paper and the attached geometric and dynamical structures. Their construction can be thought of as reverse-engineering the unit tangent bundle $T^{1}\mathbb{X}$ of a rank-one symmetric space $\mathbb{X}$ from the projections $T^{1}\mathbb{X}\to\partial_{\infty}\mathbb{X}$ sending a unit tangent vector $v\in T^{1}_{x}\mathbb{X}$ to the future and past endpoints $v^{\pm}\in\partial_{\infty}\mathbb{X}$ of the geodesic initiated by $v$. We replace the visual boundary $\partial_{\infty}\mathbb{X}$ with a pair of opposite flag manifolds ±$\mathcal{F}^{\pm}\simeq G/P^{\pm}_{\Theta}$. The image in $\partial_{\infty}{\mathbb{X}}\times\partial_{\infty}{\mathbb{X}}$ of the projection $v\mapsto(v^{+},v^{-})$ will be replaced with the transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\subset{\mathcal{F}}^{+}\times{\mathcal{F}}^{-}$, the unique open $G$-orbit in +×${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\simeq G/P^{+}_{\Theta}\times G/P^{+}_{\Theta}$ which can be equivalently described as the homogeneous space +×${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\simeq G/L_{\Theta}$.
@@ 1.1.3. $(G,X)$-structures for Anosov subgroups
Starting with a discrete subgroup $\Gamma<G$, one way of producing a $(G,X)$-manifold is to find a non-empty open $\Gamma$-invariant subset $U\subset X$ on which the $\Gamma$-action is free and properly discontinuous. The quotient manifold $\mathcal{M}_{\Gamma}:=\Gamma\backslash U$ carries a $(G,X)$-structure which we call uniformized (this is referred to as “Type $\mathrm{U}$” in [Kas18]).
@@ 1.2.1. Uniformized $(G,X)$-structures
Let $\Gamma<G$ be $\Theta$-Anosov, and let $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$. If $d_{e}\mathbf{b}\in\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$, then there is a non-empty $\Gamma\times\phi_{\mathbf{b}}^{t}$-invariant open set ~$\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$ on which $\Gamma$ acts properly discontinuously, yielding a dynamically complete uniformized $(G,{\mathbb{L}}_{\mathbf{b}})$-manifold $\mathcal{M}_{\mathbf{b}}:=\Gamma\backslash\widetilde{\mathcal{M}_{\mathbf{b}}}$ when $\Gamma$ is torsion-free.
@@
The homogeneous space $\mathbb{W}_{\Theta}=G/M_{\Theta}$ often appears in the study of Anosov subgroups in the form of a Hopf parametrization $\mathbb{W}_{\Theta}\simeq G/L_{\Theta}\times\mathfrak{a}_{\Theta}$ with a cocycle defining the $G$-action on the right hand side, similarly for ${\mathbb{L}}_{\mathbf{b}}\simeq G/L_{\Theta}\times\mathbb{R}.$ The novelty in Theorem A and Corollary B is the discovery of the domains of proper discontinuity ~$\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$, $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}.$
@@ 1.2.3. Ruelle zeta functions
Let $\lambda:G\to{\bf\mathfrak{a}}^{+}$ denote the Jordan projection with respect to a chosen positive Weyl chamber (see Section 3.7.1). Given a $\Theta$-Anosov subgroup $\Gamma<G$ and $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ as in Theorem C with $d_{e}{\mathbf{b}}$ canonically extended to a linear form on ${\bf\mathfrak{g}}$, the non-trivial primitive periodic orbits of the flow $\phi_{\mathbf{b}}^{t}$ are in bijection with primitive hyperbolic conjugacy classes in $\Gamma$ and the period of $[\gamma]$ is equal to $d_{e}{\mathbf{b}}(\lambda(\gamma)).$ Since the flow is real analytic, results of Fried [Fri95] imply the meromorphic continuation of Ruelle zeta functions twisted by linear representations, which immediately yields:
@@
coverges for $\mathrm{Re}(s)\gg 1$ and admits a meromorphic continuation to $s\in\mathbb{C}.$ Here $[\Gamma]_{\mathrm{prim}}$ is the set of conjugacy classes of non-trivial primitive elements in $\Gamma$ and the appearance of $\rho([\gamma])$ is only through its conjugation invariant characteristic polynomial.
@@
When $\Gamma<G$ is Zariski-dense, Chow-Sarkar [CS24b] recently elaborated the estimates necessary to prove a type of local non-integrability condition (uniformly in the admissible parameter $\mathbf{b}$) for Sambarino’s refraction flows, leading to a proof that the usual Ruelle zeta functions $\zeta_{\mathbf{b}}$ have a zero/pole-free strip to the left of their simple pole at the topological entropy $h_{\mathrm{top}}(\psi_{\mathbf{b}}^{t})$, thus establishing exponential mixing for the measure of maximal entropy. Strictly speaking, this result is only true when $\mathbf{b}$ satisfies the hypotheses of Theorem C.(3),.(3),. ( 3 ) , though in general one may pass to a smaller subset of roots and obtain the same result (see [CS24b, Remark $1.3$]).
Similar results were previously obtained for irreducible projective Anosov subgroups and arbitrary Gibbs measures in [DMS24], where the local non-integrability condition is geometrically manifested by the existence of a contact structure on the ambient Axiom A system. In general, Theorem 3 shows that under the failure of Theorem C.(3).(3). ( 3 ) and therefore the failure of the Axiom A system to be contact, there is a canonical contact reduction (e.g. quotient), which reveals that all refraction flows may be embedded in a contact Axiom A system, and provides a differential-geometric mechanism for the local non-integrability in consonance with the results of [CS24b]. Put simply, the success/failure of local non-integrability is exactly captured by whether the ambient Axiom A system is contact, and the fact that the contact property can be recovered via contact reduction (Theorem 3) is in harmony with [CS24b, Remark $1.3$]. Looking towards the future, exponential mixing for arbitrary Gibbs measure and with no Zariski-density/irreducibility hypothesis on $\Gamma<G$ remains an open problem; an analysis in the style of [DMS24] utilizing the work of Stoyanov [Sto11] is one promising avenue of attack.
@@ 2.1. Dynamics
[…]
@@ 3.1.2. The structure of parabolic subgroups
$1\rightarrow N_{X}\rightarrow P_{X}\rightarrow P_{X}/N_{X}\rightarrow 1$
@@ 3.2. Flag manifolds and transverse flag spaces
Consider two flag manifolds $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$. A pair $(\mathbf{x}^{+},\mathbf{x}^{-})\in\mathcal{F}^{+}\times\mathcal{F}^{-}$ is transverse if there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ such that $\mathrm{Stab}_{G}(\mathbf{x}^{+})=P_{X}$ and $\mathrm{Stab}_{G}(\mathbf{x}^{-})=P_{-X}$. We write $\mathbf{x}^{+}\pitchfork\mathbf{x}^{-}$ if $(\mathbf{x}^{+},\mathbf{x}^{-})$ is transverse and denote the set of all transverse pairs in $\mathcal{F}^{+}\times\mathcal{F}^{-}$ by ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ (or ${{\mathcal{F}}^{\pitchfork}}$ for short, when the factors ${\mathcal{F}}^{+}$, ${\mathcal{F}}^{-}$ are clear from the context).
The flag manifolds $\mathcal{F}^{+},\mathcal{F}^{-}$ are opposite if ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\neq\emptyset$. In this case, we call ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ a transverse flag space.
@@
which justifies the terminology “transverse” and shows that the set ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, which by definition is a single $G$-orbit, is open and dense in ${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}$.
Every transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ comes with a pair of projections
${{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}}$${{\mathcal{F}}^{+}}$${{\mathcal{F}}^{-}}$$\scriptstyle{p_{{\mathcal{F}}^{+}}}$$\scriptstyle{p_{{\mathcal{F}}^{-}}}$ (3.7)
given by the restriction to ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ of the canonical projections ${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\to{\mathcal{F}}^{\pm}$.
@@ 3.3. Pseudo-Riemannian structure of transverse flag spaces
The tangent space of a transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ at a point $(\mathbf{x}^{+},\mathbf{x}^{-})$ has a nice description: denoting by ${\bf\mathfrak{p}}^{+}$ and ${\bf\mathfrak{p}}^{-}$ the Lie algebras of their respective stabilizers and $\mathfrak{l}={\bf\mathfrak{p}}^{+}\cap{\bf\mathfrak{p}}^{-}$, the projection of $\mathfrak{n}_{\mathfrak{p}^{+}}\oplus\mathfrak{n}_{\mathfrak{p}^{-}}\subset{\bf\mathfrak{g}}$ onto ${\bf\mathfrak{g}}/\mathfrak{l}$ is an isomorphism, and the latter can be identified with $T_{(\mathbf{x}^{+},\mathbf{x}^{-})}{{\mathcal{F}}^{\pitchfork}}$ via the derivative of the $G$-action.
@@ 3.4. Standard parabolic subgroups and flag manifolds
Let $\theta:{\bf\mathfrak{g}}\to{\bf\mathfrak{g}}$ be a Cartan involution defining a Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ into its $+1$ and $-1$ eigenspaces, which are orthogonal with respect to the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$, and let $K<G$ be the maximal compact subgroup with Lie algebra ${\bf\mathfrak{k}}$. It is the fixed point set of the global Cartan involution, which by abuse of notation we again denote by $\theta:G\to G$. The Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ is $\mathrm{Ad}(K)$-invariant.
@@ 3.5. The hierarchy of flag manifolds
Let ${\mathcal{F}},{\mathcal{F}}^{\prime}$ be two flag manifolds. If there exists a $G$-equivariant map ${\mathcal{F}}\to{\mathcal{F}}^{\prime}$, then it is unique and it is a fiber bundle projection.
@@
Let ${\mathcal{F}},{\mathcal{F}}^{\prime}$ be two flag manifolds. We say that ${\mathcal{F}}$ dominates ${\mathcal{F}}^{\prime}$, and denote it by ${\mathcal{F}}^{\prime}\prec{\mathcal{F}}$, if there is a $G$-equivariant map $\Pi^{\mathcal{F}}_{\mathcal{F}^{\prime}}:{\mathcal{F}}\to{\mathcal{F}}^{\prime}$.
@@
Let ${\mathcal{F}}$ be a standard flag manifold. Proposition 3.14 tells us that for a fixed choice of restricted root space decomposition with closed Weyl chamber ${\bf\mathfrak{a}}^{+}$ there is $X\in{\bf\mathfrak{a}}^{+}$ such that ${\mathcal{F}}={\mathcal{F}}_{X}={\mathcal{F}}_{\Theta_{X}}$ in the notation of Definition 3.12, and the set ${\Theta_{X}}$ determines ${\mathcal{F}}$ uniquely. Now ${\mathcal{F}}_{-X}$ is opposite to ${\mathcal{F}}$, and by Fact 3.11 the adjoint $G$-orbit of $-X$ intersects $\mathfrak{a}^{+}$ in a unique element $X^{\prime}$, so that ${\mathcal{F}}_{-X}={\mathcal{F}}_{\Theta_{X^{\prime}}}$ is uniquely determined by $\Theta_{X^{\prime}}$. Concretely $X^{\prime}$ is given by $X^{\prime}=\iota_{\mathrm{o}}(X)$ and we get $\Theta_{X^{\prime}}=\iota_{\mathrm{o}}(\Theta_{X})$. In particular, $\Theta_{X^{\prime}}$ is uniquely determined by $\Theta_{X}$ If $Y\in{\bf\mathfrak{a}}^{+}$ is another element such that ${\mathcal{F}}={\mathcal{F}}_{Y}$, then by Proposition 3.14 we have $\Theta_{Y}=\Theta_{X}$ and thus ${\mathcal{F}}_{-Y}={\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}={\mathcal{F}}_{-X}$. Thus, ${\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}$ is the unique opposite of ${\mathcal{F}}_{\Theta_{X}}$ and the proof is finished. ∎
Given a standard flag manifold ${\mathcal{F}}$, we write $\overline{{\mathcal{F}}}$ for its opposite standard flag manifold. The notions of “opposition” for standard flag manifolds and parabolic subgroups are compatible in the sense that the opposite of ${\mathcal{F}}_{\Theta}=G\cdot{\bf\mathfrak{p}}_{\Theta}$ is ¯$\overline{\mathcal{F}}_{\Theta}={\mathcal{F}}_{\iota_{o}(\Theta)}=G\cdot\overline{\bf\mathfrak{p}}_{\Theta}$, where $\overline{\bf\mathfrak{p}}_{\Theta}=\theta({\bf\mathfrak{p}}_{\Theta})$ is the image under the Cartan involution of the Lie algebra of $P_{\Theta}$.

### v3

@@ 1.1. Dynamical $(G,X)$-structures
[…]
@@ 1.1.3. $(G,X)$-structures for Anosov subgroups
Starting with a discrete subgroup $\Gamma<G$, one way of producing a $(G,X)$-manifold is to find a non-empty open $\Gamma$-invariant subset $U\subset X$ on which the $\Gamma$-action is free and properly discontinuous. The quotient manifold $\U\mathcal{M}_{\Gamma}:=\Gamma\backslash$ U carries a $(G,X)$-structure which we call uniformized (this is referred to as “Type $\mathrm{U}$” in [Kas18]).
@@ 1.2.1. Uniformized $(G,X)$-structures
Let $\Gamma<G$ be $\Theta$-Anosov, and let $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$. If $d_{e}\mathbf{b}\in\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$, then there is a non-empty $\Gamma\times\phi_{\mathbf{b}}^{t}$-invariant open set $\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$ on which $\Gamma$ acts properly discontinuously, yielding a dynamically complete uniformized $(G,{\mathbb{L}}_{\mathbf{b}})$-manifold $\mathcal{M}_{\mathbf{b}}:=\Gamma\backslash\widetilde{\mathcal{M}_{\mathbf{b}}}$ when $\Gamma$ is torsion-free.
@@
The homogeneous space $\mathbb{W}_{\Theta}=G/M_{\Theta}$ often appears in the study of Anosov subgroups in the form of a Hopf parametrization $\mathbb{W}_{\Theta}\simeq G/L_{\Theta}\times\mathfrak{a}_{\Theta}$ with a cocycle defining the $G$-action on the right hand side, similarly for ${\mathbb{L}}_{\mathbf{b}}\simeq G/L_{\Theta}\times\mathbb{R}.$ The novelty in Theorem A and Corollary B is the discovery of the domains of proper discontinuity $\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$, $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}.$
@@ 1.2.3. Ruelle zeta functions
Let $\lambda:G\to{\bf\mathfrak{a}}^{+}$ denote the Jordan projection with respect to a chosen positive Weyl chamber (see Section 3.7.1). Given a $\Theta$-Anosov subgroup $\Gamma<G$ and $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ as in Theorem C with $d_{e}{\mathbf{b}}$ canonically extended to a linear form on $\bf\mathfrak{g}$}, the non-trivial primitive periodic orbits of the flow $\phi_{\mathbf{b}}^{t}$ are in bijection with primitive hyperbolic conjugacy classes in $\Gamma$ and the period of $\gamma]$ is equal to $d_{e}{\mathbf{b}}(\lambda(\gamma)).$ Since the flow is real analytic, results of Fried [Fri95] imply the meromorphic continuation of Ruelle zeta functions twisted by linear representations, which immediately yields:
@@
converges for $\mathrm{Re}(s)\gg 1$ and admits a meromorphic continuation to $s\in\mathbb{C}.$ Here $\Gamma]_{\mathrm{prim}}$ is the set of conjugacy classes of non-trivial primitive elements in $\Gamma$ and the appearance of $\rho([\gamma])$ is only through its conjugation invariant characteristic polynomial.
@@
When $\Gamma<G$ is Zariski-dense, Chow-Sarkar [CS24b] recently elaborated the estimates necessary to prove a type of local non-integrability condition (uniformly in the admissible parameter $\mathbf{b}$) for Sambarino’s refraction flows, leading to a proof that the usual Ruelle zeta functions $\zeta_{\mathbf{b}}$ have a zero/pole-free strip to the left of their simple pole at the topological entropy $h_{\mathrm{top}}(\psi_{\mathbf{b}}^{t})$, thus establishing exponential mixing for the measure of maximal entropy. Strictly speaking, this result is only true when $\mathbf{b}$ satisfies the hypotheses of Theorem C.(3),.(3), though in general one may pass to a smaller subset of roots and obtain the same result (see [CS24b, Remark $1.3$]).
Similar results were previously obtained for irreducible projective Anosov subgroups and arbitrary Gibbs measures in [DMS24], where the local non-integrability condition is geometrically manifested by the existence of a contact structure on the ambient Axiom A system. In general, Theorem 3 shows that under the failure of Theorem C.(3).(3) and therefore the failure of the Axiom A system to be contact, there is a canonical contact reduction (e.g. quotient), which reveals that all refraction flows may be embedded in a contact Axiom A system, and provides a differential-geometric mechanism for the local non-integrability in consonance with the results of [CS24b]. Put simply, the success/failure of local non-integrability is exactly captured by whether the ambient Axiom A system is contact, and the fact that the contact property can be recovered via contact reduction (Theorem 3) is in harmony with [CS24b, Remark $1.3$]. Looking towards the future, exponential mixing for arbitrary Gibbs measure and with no Zariski-density/irreducibility hypothesis on $\Gamma<G$ remains an open problem; an analysis in the style of [DMS24] utilizing the work of Stoyanov [Sto11] is one promising avenue of attack.
1.3. Anosov subgroups vs. Anosov flows
Hyperbolic homogeneous flows play an important role in the proof of a famous result of Benoist-Foulon-Labourie [BFL92]: if an Anosov flow $\varphi^{t}:M\to M$ preserves a contact form on a compact manifold $M$ and has smooth stable and unstable distributions, it is smoothly conjugate to a reparametrisation of the geodesic flow of a closed locally symmetric space of rank one. Their proof can be broken down into the following steps: first proving that the flow $\varphi^{t}$ is induced from a $(G,\mathbb{L}_{\mathbf{b}})$-structure for some semi-simple Lie group $G$, then proving that this structure is complete (i.e. $M$ is covered by $\mathbb{L}_{\mathbf{b}}$), and finally that $G$ has rank one. In particular, the locally homogeneous flow $(\mathcal{M}_{\mathbf{b}},\phi^{t}_{\mathbf{b}})$ that we construct from an Anosov subgroup cannot be an Anosov flow if $G$ is simple and has higher rank.
@@ 2.1. Dynamics
Let $\mathcal{K}\subset{\mathcal{M}}$ be a compact $\phi^{t}$-invariant set and $E$ a continuous vector bundle over $\mathcal{K}$ equipped with a continuous flow $\phi_{E}^{t}:E\to E$ lifting $\phi^{t}$ over $\mathcal{K}$. Then $\phi_{E}^{t}$ is uniformly contracting (resp. expanding) on $E$ if for some (hence any) continuous bundle norm $\left\|\cdot\right\|$ on $E$ there are constants $C,c>0$ such that for all $p\in\mathcal{K}$ and all $v\in E_{p}$ one has
@@ 3.1.2. The structure of parabolic subgroups
$1\rightarrow N_{X}\rightarrow P_{X}\rightarrow P_{X}/N_{X}\rightarrow 1.$
@@ 3.2. Flag manifolds and transverse flag spaces
[…]
@@
[…]
@@ 3.3. Pseudo-Riemannian structure of transverse flag spaces
[…]
@@ 3.4. Standard parabolic subgroups and flag manifolds
Let $\theta:{\bf\mathfrak{g}}\to{\bf\mathfrak{g}}$ be a Cartan involution defining a Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ into its +1+1 and 1-1 eigenspaces, which are orthogonal with respect to the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$, and let $K<G$ be the maximal compact subgroup with Lie algebra $\bf\mathfrak{k}$}. It is the fixed point set of the global Cartan involution, which by abuse of notation we again denote by $\theta:G\to G$. The Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ is $\mathrm{Ad}(K)$-invariant.
@@ 3.5. The hierarchy of flag manifolds
Let $\mathcal{F}$$\mathcal{F}$$^{\prime}$ be two flag manifolds. If there exists a $G$-equivariant map $\mathcal{F}$$\to{\mathcal{F}}^{\prime}$, then it is unique and it is a fiber bundle projection.
@@
Let $\mathcal{F}$$\mathcal{F}$$^{\prime}$ be two flag manifolds. We say that $\mathcal{F}$} dominates $\mathcal{F}$$^{\prime}$, and denote it by $\mathcal{F}$$^{\prime}\prec{\mathcal{F}}$, if there is a $G$-equivariant map $\Pi^{\mathcal{F}}_{\mathcal{F}^{\prime}}:{\mathcal{F}}\to{\mathcal{F}}^{\prime}$.
@@
Let $\mathcal{F}$} be a standard flag manifold. Proposition 3.14 tells us that for a fixed choice of restricted root space decomposition with closed Weyl chamber $\bf\mathfrak{a}$$^{+}$ there is $X\in{\bf\mathfrak{a}}^{+}$ such that ${\mathcal{F}}={\mathcal{F}}_{X}={\mathcal{F}}_{\Theta_{X}}$ in the notation of Definition 3.12, and the set $\Theta_{X}$} determines $\mathcal{F}$} uniquely. Now ${\mathcal{F}}_{-X}$ is opposite to $\mathcal{F}$}, and by Fact 3.11 the adjoint $G$-orbit of X-X intersects $\mathfrak{a}^{+}$ in a unique element $X^{\prime}$, so that ${\mathcal{F}}_{-X}={\mathcal{F}}_{\Theta_{X^{\prime}}}$ is uniquely determined by $\Theta_{X^{\prime}}$. Concretely $X^{\prime}$ is given by $X^{\prime}=\iota_{\mathrm{o}}(X)$ and we get $\Theta_{X^{\prime}}=\iota_{\mathrm{o}}(\Theta_{X})$. In particular, $\Theta_{X^{\prime}}$ is uniquely determined by $\Theta_{X}$ . If $Y\in{\bf\mathfrak{a}}^{+}$ is another element such that ${\mathcal{F}}={\mathcal{F}}_{Y}$, then by Proposition 3.14 we have $\Theta_{Y}=\Theta_{X}$ and thus ${\mathcal{F}}_{-Y}={\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}={\mathcal{F}}_{-X}$. Thus, ${\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}$ is the unique opposite of ${\mathcal{F}}_{\Theta_{X}}$ and the proof is finished. ∎
Given a standard flag manifold $\mathcal{F}$}, we write $\overline{{\mathcal{F}}}$ for its opposite standard flag manifold. The notions of “opposition” for standard flag manifolds and parabolic subgroups are compatible in the sense that the opposite of ${\mathcal{F}}_{\Theta}=G\cdot{\bf\mathfrak{p}}_{\Theta}$ is $\overline{{\mathcal{F}}}_{\Theta}={\mathcal{F}}_{\iota_{o}(\Theta)}=G\cdot\overline{{\bf\mathfrak{p}}}_{\Theta}$, where $\overline{{\bf\mathfrak{p}}}_{\Theta}=\theta({\bf\mathfrak{p}}_{\Theta})$ is the image under the Cartan involution of the Lie algebra of $P_{\Theta}$.